# Open-TeleVision: Teleoperation with Immersive Active Visual Feedback

### v1

@@ 1 Introduction
Learning-based robotic manipulation has advanced to a new level in the past few years, thanks to large-scale real-robot data [1, 2]. Teleoperation has been playing an important role in data collection for imitation learning, where it not only offers accurate and precise manipulation demonstrations, but also provides natural and smooth trajectories that allow the learned policies to generalize to new environment configurations and tasks. Various teleoperation approaches have been studied using VR devices [3, 2], RGB cameras [4, 5, 6], wearable gloves [7, 8, 9], and customized hardwares [10, 11].
@@
Our experiments follow the teleoperation and imitation learning paradigm. We experiment with multiple fine-grained manipulation tasks across two robots as shown in Fig. LABEL:fig:teaser. For teleoperation, we qualitatively show the benefits of using an active camera allows more intuitive and focused observation, and quantitatively show streaming stereoscopic video allows better success rate and shorter completion time across different users. For imitation learning, we quantitatively report active camera sensing enables faster inference for real-time smooth control, and using stereoscopic inputs achieves better manipulation performance. Our policy can conduct long-horizon tasks such as inserting multiple cans in a sequence. A key benefit of our system is enabling remote control by an operator via the Internet. One of the authors, Ge Yang at MIT (east coast) is able to teleoperate the H1 robot at UC San Diego (west coast). We include videos of this type of coast-to-coast teleoperations on our website.
@@ 3.1 Imitation Learning Results
We ablate our key design choices of the system and show from real-world experiments their effectiveness. The two baselines are w. ResNet18, which uses the original ACT visual backbone ResNet18 [31], and w/o Stereo Input, which only takes the visual tokens from the left image instead of both. All models are trained using AdamW optimizer [32, 33] with a learning rate of $55e-$ 5, a batch size of $45$ and for $25k$ iterations on a single RTX 4090 GPU. We conduct Can Sorting task on both H1 and GR-1 robots. All other tasks are conducted with H1 robot only.
From our experiment results shown in Tab. 1, we notice that the original ResNet backbone fails to adequately perform all four tasks.
In the original ACT implementation, four cameras are used (two fixed, two wrist mounted) to alleviate the deficiency of explicit spatial information from using RGB images. However, our setup involves only two stereo RGB cameras, potentially making spatial information retrieval more difficult for ResNet backbone.
@@ 4 Related Work
Imitation learning. The topic of imitation learning has covered a wide range of literature. If we distinguish the existing work with sources of demonstrations, we can classify them into learning from real-robot expert data [14, 10, 12, 59, 60, 61, 62, 50, 63, 1, 64, 51, 65], learning from play data [66, 67, 68], and learning from human demonstrations [67, 69, 70, 71, 72, 73, 74, 56]. Besides learning for manipulation, motion imitation for physical characters and real robots have also been widely studied in  [75, 76, 40, 77, 78, 79, 80, 81, 82, 80, 83, 84, 85, 86, 87, 88]. This paper falls into imitation learning for manipulation tasks using real-robot data collected by human teleoperation.

### v2

@@ 1 Introduction
Learning-based robotic manipulation has advanced to a new level in the past few years, thanks to large-scale real-robot data [1, 2]. Teleoperation has been playing an important role in data collection for imitation learning, where it not only offers accurate and precise manipulation demonstrations, but also provides natural and smooth trajectories that allow the learned policies to generalize to new environment configurations and tasks. Various teleoperation approaches have been studied using VR devices [3, 2, 4, 5], RGB cameras [6, 7, 8], wearable gloves [9, 10, 11], and customized hardwares [12, 13].
@@
Our experiments follow the teleoperation and imitation learning paradigm. We experiment with multiple fine-grained manipulation tasks across two robots as shown in Fig. LABEL:fig:teaser. For teleoperation, we qualitatively show the benefits of using an active camera allows more intuitive and focused observation, and quantitatively show streaming stereoscopic video allows better success rate and shorter completion time across different users. For imitation learning, we quantitatively report active camera sensing enables faster inference for real-time smooth control, and using stereoscopic inputs achieves better manipulation performance. Our policy can conduct long-horizon tasks such as inserting multiple cans in a sequence. A key benefit of our system is enabling remote control by an operator via the Internet. One of the authors, Ge Yang at MIT (east coast) is able to teleoperate the H1 robot at UC San Diego (west coast). We include coast-to-coast teleoperation videos on our website.
@@ 3.1 Imitation Learning Results
We ablate our key design choices of the system and show from real-world experiments their effectiveness. The two baselines are w. ResNet18, which uses the original ACT visual backbone ImageNet Pre-trained ResNet18 [33], and w/o Stereo Input, which only takes the visual tokens from the left image instead of both. All models are trained using AdamW optimizer [34, 35] with a learning rate of $55e-$ 5, a batch size of $45$ and for $25k$ iterations on a single RTX 4090 GPU. We conduct Can Sorting task on both H1 and GR-1 robots. All other tasks are conducted with H1 robot only.
From our experiment results shown in Tab. 1, we notice that the original ResNet backbone (ImageNet pre-trained) fails to adequately perform all four tasks. In the original ACT implementation, four cameras are used (two fixed, two wrist-mounted) to alleviate the deficiency of explicit spatial information from using RGB images. Our setup involves only two stereo RGB cameras, potentially making spatial information retrieval more challenging for ResNet backbone.
@@ 4 Related Work
Imitation learning. The topic of imitation learning has covered a wide range of literature. If we distinguish the existing work with sources of demonstrations, we can classify them into learning from real-robot expert data [16, 12, 14, 61, 62, 63, 64, 52, 65, 1, 66, 53, 67, 68], learning from play data [69, 70, 71], and learning from human demonstrations [70, 72, 73, 74, 75, 76, 77, 58]. Besides learning for manipulation, motion imitation for physical characters and real robots have also been widely studied in  [78, 79, 42, 80, 81, 82, 83, 84, 85, 83, 86, 87, 88, 89, 90, 91]. This paper falls into imitation learning for manipulation tasks using real-robot data collected by human teleoperation.